OpenStudy (kamille):

Hello again! I know that: f(x)=3/x and g(x)=x-1 And need to find f(g(2)) i think I need to do like that: \[\frac{ 3 }{ x(x-1) }=\frac{ 3 }{ x ^{2}-x }\] and then insert 2 where is x, is it right or not?

OpenStudy (mertsj):

The easiest way is to replace x with 2 in the function g. Then take that answer and replace x in the function f with it. That will be the f(g(2))

OpenStudy (kamille):

fak, I thoughet I made a mistake. then realized I havent :/ so I get: g(2)=2-1=1 f(g(2))=3/1=3?

OpenStudy (mertsj):

That's what I got.

OpenStudy (kamille):

thank you! I have helped me a lot recently: )

OpenStudy (mertsj):

Glad to be of service.

OpenStudy (kamille):

you* haha I think I am too tired :/